# StreamSplat: Towards Online Dynamic 3D Reconstruction from Uncalibrated Video Streams

### v1

@@ Appendix B Implementation Details
Model Configurations. As shown in Table 5, StreamSplat consists of an image tokenizer, a static encoder, a dynamic decoder, and an upsampler. The image tokenizer takes RGBD inputs at a resolution of 288$\times$512 and uses a patch size of 8 to produce 768-dimensional tokens. Both the static encoder and the dynamic decoder contain 10 transformer layers with an embedding dimension of 768 and 12 attention heads. The dynamic decoder uses 0.1 stochastic drop path to prevent overfitting on image features. An upsampler with 2 layers of window attention with 2304 token length increases the token length to 16×16$\times$16 × and reduces the embedding dimension from 768 to 192. The model uses a linear head for static prediction and a two-layer MLP for deformation prediction. We apply loss balancing weights of 1.0 for MSE, 0.05 for both LPIPS and depth, and 3.0 for the mask loss.
3DGS Parameterization. 3DGS parameters are obtained as summarized in Algorithm 1. A linear head projects each Gaussian token $\hat{\mathbf{E}}$ and, through parameter-specific activations, yields the static tuple $(\boldsymbol{\mu}_{0},\mathbf{S},\mathbf{R},\boldsymbol{\alpha}_{0},\boldsymbol{c})$. A 2-layer MLP, conditioned on the concatenation $\hat{\mathbf{E}}\oplus\hat{\mathbf{d}}$, predicts the velocity $\mathbf{v}$ and opacity coefficients $\boldsymbol{\gamma}$, from which the final time-dependent values are computed by integrating the current time $t$. Note, we adopt our proposed Probabilistic Position Sampling for all position-related parameters, including position offset $\boldsymbol{o}$ and velocity $\mathbf{v}$, to improve robustness and avoid spatial local minima.
Figure 8: Qualitative results with 8-frame interval on different datasets. Blue box: given frames; Red box: interpolated frames.
Figure 9: Ablation. w/o bi.: without bidirectional deformation field. Blue box: given frames; Red box: interpolated frames.
@@ Appendix C Additional Experimental Results
Due to the limited space in the main manuscript, we provide additional experimental results in this section, including qualitative comparisons with other dynamic methods on static scene reconstruction (Figure 7), and qualitative comparisons on ablation studies of our bidirectional deformation field (Figure 9), and more qualitative results on video reconstruction and interpolation on CO3Dv2 [31], DAVIS [33], and Youtube-VOS [34] (Figure 8). Please refer to the website for video results.
Appendix D Limitations and Impact
@@
While StreamSplat achieves strong performance, certain design choices introduce some limitations. First, the framework relies on pseudo-depth maps predicted by an external monocular estimator, which may introduce noise–particularly around fine-scale geometry and depth discontinuities. To mitigate this, we incorporate an adaptive decay weighting scheme during training to downweight unreliable depth supervision. Nonetheless, scaling the training dataset to enable internal depth refinement remains a promising direction to reduce reliance on external priors. Second, the bidirectional deformation field is trained over a two-frame window for efficiency. As a result, information from earlier frames may be lost in dynamic scenes with fast motion or extended occlusions. Future work may explore efficient mechanisms for adaptively selecting and fusing Gaussians across extended frame histories, which could help retain more temporal context in challenging scenarios.
Societal impact. While our work enables instant dynamic 3D perception from monocular video, potentially improving accessibility and safety real-world applications like autonomous driving systems, it also warrants careful evaluation in edge cases where performance limitations could affect reliability in safety-critical scenarios.
Appendix E Licenses
@@
  • CO3Dv2 [31]: CC BY-NC 4.0
  • RealEstate10K [32]: CC BY 4.0
  • DAVIS [33]: BSD 3-Clause License
  • YouTube-VOS [34]: CC BY 4.0
@@
  • DepthAnythingv2 [65]: Apache-2.0 License and CC BY-NC 4.0

### v2

@@ Appendix B Implementation Details
[…]
3DGS Parameterization. 3DGS parameters are obtained as summarized in Algorithm 1. A linear head projects each Gaussian token $\hat{\mathbf{E}}$ and, through parameter-specific activations, yields the static tuple $(\bm{\mu}_{0},\mathbf{S},\mathbf{R},\bm{\alpha}_{0},\bm{c})$. A 2-layer MLP, conditioned on the concatenation $\hat{\mathbf{E}}\oplus\hat{\mathbf{d}}$, predicts the velocity $\mathbf{v}$ and opacity coefficients $\bm{\gamma}$, from which the final time-dependent values are computed by integrating the current time $t$. Note, we adopt our proposed Probabilistic Position Sampling for all position-related parameters, including position offset $\bm{o}$ and velocity $\mathbf{v}$, to improve robustness and avoid spatial local minima. Spatial scaling factors $(f_{x},f_{y})=(256,144)$, which equals to patch numbers.
Dynamic Deformation Decoding. As shown in Figure 9, given two consecutive frames from a randomly sampled time interval, $I_{t_{1}}$ and $I_{t_{2}}$, we first use the frozen static encoder to extract their 3DGS Embeddings $\mathbf{h}_{t_{1}}$ and $\mathbf{h}_{t_{2}}$. We also extract DINOv2 features Oquab et al. (2023) $\mathbf{f}_{t_{1}}$ and $\mathbf{f}_{t_{2}}$ from $I_{t_{1}}$ and $I_{t_{2}}$, respectively. To distinguish the embeddings, we add learnable Type Embeddings Devlin et al. (2019); Chen et al. (2020) $\mathbf{T}_{src},\mathbf{T}_{tgt}\in\mathbb{R}^{D}$ to $\mathbf{h}_{t_{1}}$ and $\mathbf{h}_{t_{2}}$, yielding $\hat{\mathbf{h}}_{t_{1}}$ and $\hat{\mathbf{h}}_{t_{2}}$. These are processed by the decoder as follows:
$=\text{Self-Attn}([\hat{\mathbf{h}}_{t_{1}},\hat{\mathbf{h}}_{t_{2}}]),$ (11)
$\hat{\mathbf{h}}_{t_{1}}$ $=\text{Cross-Attn}(\hat{\mathbf{h}}_{t_{1}},\mathbf{f}_{t_{2}}),\ \hat{\mathbf{h}}_{t_{2}}=\text{Cross-Attn}(\hat{\mathbf{h}}_{t_{2}},\mathbf{f}_{t_{1}})$
where $\cdot$] denotes concatenation. After passing through a few decoder blocks, we obtain the Deformation Embeddings $\mathbf{d}_{t_{1}},\mathbf{d}_{t_{2}}]\in\mathbb{R}^{2N\times D}$, which are then upsampled via the same upsampler to produce deformation tokens $\hat{\mathbf{d}}\in\mathbb{R}^{32N\times D/4}$. Each deformation token $\hat{\mathbf{d}}_{j}$ is concatenated with the corresponding static Gaussian token from the same frame $\hat{\mathbf{E}}_{j}$ to form a joint token $\hat{\mathbf{E}}_{j}\oplus\hat{\mathbf{d}}_{j}$, which is passed through a 2-layer MLP head to predict the deformation field, including velocity $\mathbf{v}_{j}$ and opacity coefficient $\bm{\gamma}_{j}$. These deformation fields allow Gaussians to move and fade over time, enabling computation of dynamic 3D Gaussians at arbitrary times for continuous scene reconstruction.
Figure 10: Qualitative results on RE10K. StreamSplat produces detailed and consistent reconstructions across diverse scenes.
Method Rep. Type Given View Novel View Average
PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS
pixelSplat Charatan et al. (2024) Stat. 30.70 0.952 0.055 28.31 0.905 0.097 28.99 0.918 0.085
MVSplat Chen et al. (2024) 31.48 0.962 0.046 28.48 0.909 0.091 29.34 0.924 0.078
NoPoSplat Ye et al. (2024) 29.50 0.939 0.069 28.65 0.913 0.096 28.90 0.920 0.089
CoDeF Ouyang et al. (2024) Dyn. 35.13 0.943 0.091 20.51 0.591 0.402 21.77 0.625 0.374
DGMarbles Stearns et al. (2024) 27.48 0.867 0.232 23.40 0.727 0.333 23.73 0.738 0.325
Figure 11: Qualitative results on RE10K. Blue box: given frames; Red box: novel frames. Compared to other dynamic-scene reconstruction methods, StreamSplat produces more detailed and consistent 3D reconstructions across diverse scenes, whereas other methods often exhibit distortions in both color and geometry.
@@ Appendix C Additional Experimental Results
Due to the limited space in the main manuscript, we provide additional experimental results in this section, including qualitative and quantitative comparisons with other methods on static scene reconstruction (Figure 10 and 11, Table 5), qualitative comparisons on ablation studies of our bidirectional deformation field (Figure 13), more qualitative results on video reconstruction on CO3Dv2 Reizenstein et al. (2021), DAVIS Pont-Tuset et al. (2017), and Youtube-VOS Xu et al. (2018) (Figure 12), and zero-shot evaluation on DyCheck Gao et al. (2022) and NVIDIA Dynamic Scene Yoon et al. (2020) benchmarks (Figure 14, Tables 6 and 7). Please refer to the project website for video results.
Table 6: DyCheck results. We report PSNR/LPIPS on novel view synthesis.
Method Extr. Intr. Apple Block Spin Windmill Space-out Teddy Wheel Average Time
4DGS (w/ cam) $\checkmark$ $\checkmark$ 15.41 / .45 11.28 / .63 14.41 / .34 15.60 / .30 14.60 / .37 12.36 / .47 11.79 / .44 13.64 / .43 >6h
DGMarbles (w/ cam) $\checkmark$ $\checkmark$ 17.70 / .49 17.42 / .38 18.88 / .43 17.04 / .39 15.94 / .44 13.95 / .44 16.14 / .35 16.72 / .42 >5h
4DGS (w/o pose) $\times$ $\checkmark$ 14.44 / .72 12.30 / .71 12.77 / .70 14.46 / .79 14.93 / .64 11.86 / .73 10.99 / .80 13.11 / .73 >6h
DGMarbles (w/o pose) $\times$ $\checkmark$ 16.50 / .50 16.11 / .36 17.51 / .42 16.19 / .45 15.97 / .44 13.68 / .44 14.58 / .39 15.79 / .43 >5h
DGMarbles (w/o cam) $\times$ $\times$ 9.91 / .85 10.21 / .81 10.64 / .74 10.32 / .71 10.82 / .71 8.33 / .84 8.10 / .76 9.76 / .77 >5h
Ours (w/o cam) $\times$ $\times$ 12.75 / .74 12.06 / .71 12.89 / .68 13.46 / .67 13.95 / .67 11.23 / .71 10.23 / .69 12.37 / .69 14s
Table 7: NVIDIA Dynamic Scenes results. We report PSNR/LPIPS on novel view synthesis.
Method Extr. Intr. Scene Info. Balloon1 Balloon2 Jumping Playground Skating Truch Umbrella Average
4DGS (w/ cam) GT GT $\checkmark$ 14.11 / .40 18.56 / .24 17.32 / .33 13.51 / .34 19.41 / .22 21.25 / .17 19.00 / .34 17.59 / .29
DGMarbles (w/ cam) GT GT $\checkmark$ 23.65 / .07 21.60 / .14 19.61 / .18 16.21 / .24 24.24 / .09 27.18 / .06 23.76 / .12 22.32 / .13
DGMarbles (w/ CUT3R) GT CUT3R $\checkmark$ 12.66 / .65 13.50 / .60 14.90 / .53 8.12 / .62 17.30 / .49 15.38 / .55 15.50 / .70 13.91 / .59
Ours (w/o cam) $\times$ $\times$ $\times$ 14.94 / .48 15.77 / .50 16.65 / .37 12.75 / .58 19.08 / .47 18.97 / .37 15.97 / .60 16.30 / .48
Static Scene Reconstruction We evaluate StreamSplat on the RE10K benchmark and compare it with recent static and dynamic reconstruction methods. For dynamic methods without camera pose input, novel views are rendered using relative timestamps. Quantitative and qualitative results are presented in Table 5 and Figures 7, 10, and 11. StreamSplat significantly outperforms all static baselines on the given-view reconstruction task. However, on novel-view reconstruction, static methods still lead, although StreamSplat achieves the best performance among all dynamic approaches.
According to our qualitative comparison in Figures 10 and 11, we attribute this to the absence of accurate camera poses input in our model. Our StreamSplat assumes smooth camera motion between two input views and relies solely on relative timestamps for novel view synthesis, which can cause slight misalignments. Furthermore, unlike static models that assume a fixed scene, StreamSplat is designed to model scene dynamics, which may introduce motion artifacts that hinder performance under static benchmarks. Despite these challenges, our StreamSplat achieves competitive average performance across all test views and consistently surpasses all dynamic baselines in every evaluation setting. Notably, StreamSplat is the only scalable approach among these that requires neither per-scene optimization Ouyang et al. (2024); Stearns et al. (2024) nor per-dataset training Charatan et al. (2024); Chen et al. (2024); Ye et al. (2024).
Figure 12: Qualitative results with 8-frame interval on different datasets. Blue box: given frames; Red box: novel frames.
Novel view synthesis on DyCheck benchmark. We have conducted a zero-shot evaluation on 7 scenes from the DyCheck iPhone dataset. To ensure fairness under uncalibrated input, we follow these settings:
  • No extrinsics: Fix camera pose and absorb motion into 3D Gaussians.
  • No intrinsics: Use a fixed camera intrinsic matrix for all methods. Specifically, for DGMarbles (w/o cam), we use GT principal point and apply only a small perturbation to the focal length to mimic no intrinsics scenarios.
  • Evaluation: Compute relative camera pose w.r.t. the fixed pose along DyCheck evaluation trajectories.
[…]
Novel view synthesis on NVIDIA Dynamic Scene benchmark. We also conducted a zero-shot evaluation on 7 scenes from NVIDIA Dynamic Scene. To ensure fairness under uncalibrated input, we evaluated DGMarbles using intrinsics estimated by CUT3R Wang et al. (2025b), while providing it with GT poses and scene metadata (scale, near/far) to ensure valid execution. The results are provided in Table 7 and Figure 14. StreamSplat (w/o cam) significantly outperforms DGMarbles (w/ CUT3R) and approaches the performance of 4DGS (w/ cam), which use both GT intrinsic and extrinsic.
Figure 13: Ablation. w/o bi.: without bidirectional deformation field. Blue box: given frames; Red box: novel frames.
Figure 14: Novel view synthesis results on DyCheck iPhone and NVIDIA Dynamic Scenes.
Figure 15: Multiple persistent Gaussians tracking across frames.
Figure 16: Qualitative results on DAVIS with frequent surface/topology changes and multi-object occlusion.
Highly dynamic scenes with topology changes. Figure 16 illustrates some representative DAVIS sequences with strong non-rigid motion and frequent topology changes, where we visualize our reconstructions across the sequence. Across dancing and acrobatic motions with large limb articulation and self-occlusion, a pedestrian entering and exiting the field of view while passing behind thin structures, and a group of people moving around and mutually occluding each other, StreamSplat preserves sharp appearance and consistent geometry over time. Despite objects appearing, disappearing, and undergoing rapid pose changes, the reconstructed frames remain temporally coherent, indicating that the bidirectional deformation and adaptive Gaussian fusion effectively handle highly dynamic, topology-changing real-world scenes.
Table 8: Runtime breakdown of StreamSplat on a single A100 GPU.
Module Time (ms) Percentage (%)
Encoder 18.8 38.4
Decoder 18.4 37.6
Rasterization 4.0 8.1
Pre-/Post-processing 7.8 15.9
Total 49 100.0
Runtime breakdown. We report a detailed runtime breakdown of StreamSplat in Table 8. On a single NVIDIA A100 GPU, the end-to-end runtime is 0.049s per frame at our default evaluation resolution. The runtime breakdown confirms the practical feasibility of near real-time online reconstruction and suggests that future speedups will primarily come from further optimizing the encoder/decoder components.
Appendix D Discussions
Our work presents StreamSplat, a fully feed-forward framework for online dynamic 3D reconstruction from uncalibrated video streams. Through extensive experiments and analysis, we have demonstrated its effectiveness while also identifying important design choices and trade-offs that need further discussion.
Orthographic Projection vs. Perspective Projection. We adopt orthographic projection as a practical choice for handling uncalibrated videos with diverse camera characteristics. This design enables StreamSplat to process videos with unknown intrinsics—ranging from standard to wide-angle and fisheye lenses—within a single model. The orthographic formulation decouples geometric reconstruction from camera parameters, absorbing camera motion into the learned Gaussian dynamics.
Our DyCheck experiments validate this choice: DGMarbles (w/o cam) achieves only 10.61 PSNR compared to our 13.46 PSNR on Windmill, demonstrating that orthographic projection doesn’t compromise accuracy when intergrate with proper deformations. While foreshortening effects occur at close range, our dynamic decoder can deform Gaussians to mimic this perspective effects, as evidenced by strong performance across RE10K and DAVIS benchmarks (Section 4). This robustness to varying camera conditions makes orthographic projection particularly suitable for real-world deployment where calibration is impractical.
In contrast, 3DGS reconstruction is highly sensitive to the assumed camera model: errors in perspective intrinsics (e.g., focal length, principal point) induce scale/shape distortions, for which we are not aware of a simple, proven remedy in online settings. Empirically, even perturbing the intrinsics of DGMarbles (w/o cam) by 10% causes a significant degradation in performance.
Opacity Deformation. The bidirectional deformation field employs time-dependent opacity to handle dynamic scene content naturally. This mechanism handles appearing and disappearing objects without explicit tracking and merging, which is crucial for online processing where scene content evolves continuously. Our ablation studies in Figure 8 reveal this: constant opacity cannot accommodate emerging or vanishing content/surfaces, resulting in ghosting and blurring effects.
Bounded Velocity. Although naturally velocity can be unbounded, we constrain it within $1,1]^{3}$ in our implementation. This choice is motivated by two considerations: 1) a velocity magnitude > 1 implies no spatial overlap between consecutive frames, which is essentially impossible in a reconstruction setting. Specifically, magnitude = 1 already means a pixel moves from one image boundary to the opposite boundary within a single frame interval, which is already a very extreme case with large motion. 2) Inferring motion for non-overlapping frames shifts the task toward video/4D generation, which is outside our scope. Thus, there is no need to allow velocity beyond magnitude 1.
Reconstruction vs. Generation. We would like to emphasize that StreamSplat performs reconstruction, not generation, distinguished by three characteristics:(1) It produces observation-constrained representations that minimize reconstruction error while maintaining temporal consistency through Adaptive Gaussian Fusion. (2) The deformation process is deterministic, and novel time/view result from geometric transformations of a fixed set of Gaussians, not probabilistic generating new Gaussians. (3) Gaussians persist across frames via soft matching in canonical space, with opacity deformation handling visibility changes rather than creating new content. Testing with large frame intervals (5-8 frames) challenges temporal coherence more rigorously than dense sampling. Our strong performance under these conditions validates the reconstruction quality, demonstrating that StreamSplat maintains scene structure even with significant temporal gaps, which is a capability essential for practical deployment where frame rates may vary.
Appendix E Limitations
@@
While StreamSplat achieves strong performance, certain design choices introduce some limitations. First, the framework relies on pseudo-depth maps predicted by an external monocular estimator, which may introduce noise–particularly around fine-scale geometry and depth discontinuities. To mitigate this, we incorporate an adaptive decay weighting scheme during training to downweight unreliable depth supervision. Nonetheless, scaling the training dataset to enable internal depth refinement remains a promising direction to reduce reliance on external priors. Second, the bidirectional deformation field is trained over a two-frame window for efficiency. As a result, information from earlier frames may be lost in dynamic scenes with fast motion or extended occlusions. Third, while orthographic projection enables robustness to uncalibrated inputs, it may introduce camera model misalignment in close-range scenes with strong perspective effects. Although our experiments show that the deformation field can largely compensate for these effects, residual distortions can remain; incorporating lightweight intrinsic estimation or perspective-aware refinement is a promising direction. Future work may involve camera estimation model to mitigate camera misalignment and explore efficient mechanisms for adaptively selecting and fusing Gaussians across extended frame histories, which could help retain more temporal context in challenging scenarios.
Appendix F Licenses
@@
  • CO3Dv2 Reizenstein et al. (2021): $C$ BY-NC 4.0
  • RealEstate10K Zhou et al. (2018): $C$ BY 4.0
  • DAVIS Pont-Tuset et al. (2017): BSD 3-Clause License
  • YouTube-VOS Xu et al. (2018): $C$ BY 4.0
@@
  • DepthAnythingv2 Yang et al. (2024b): Apache-2.0 License and $C$ BY-NC 4.0
  • DINOv2 Oquab et al. (2023): Apache-2.0 License and $C$ BY 4.0